(6 pts.] An LR circuit consists of a 1.5 kΩ resistance, 75 mH inductance, 6.0 volt battery, and a switch S. Switch S is closed at t 0, At what time will the current increase to two-thirds of its final value? 8
Answer #1

According to the given problem,

In a LR circut,

current through lr series circuit

2/3I0 = I0[1 - e-1.5*10^3*t/75*10^-3]

t = 5.493*10-5 s

t = 5.5μs
Answer #1

According to the given problem,

In a LR circut,

current through lr series circuit

2/3I0 = I0[1 - e-1.5*10^3*t/75*10^-3]

t = 5.493*10-5 s

t = 5.5μs